OpenStudy (anonymous):

e^(2x-2)-e^(x-1)-2=0

OpenStudy (anonymous):

i need the way to solve this equation

OpenStudy (anonymous):

Well ok. you can use substitution. the first term can be rewritten as e^(x-1)(2). so if we make the subst. y=e^(x-1) the above formula can be simplified to y^2 - y - 2=0... so (y-2)(y+1)=0, therefore y=2 or y=-1. now since y=e^(x-1), we have two equations.. 2=e^(x-1) and -1=e^(x-1). now from the first equation we can find the natural log of both sides and end up with ln(2)=x-1 therefore x=1+ln(2)...the second equation is invalid since ln(-1) is not defined. so ur ansa is x=1+ln(2)